How long does it take for an echo to return to a woman standing 400 ft from a cliff? Note: The velocity of sound in air is about 1100 ft/s.



1
Expert's answer
2021-09-20T14:42:11-0400
"2d=vt,""t=\dfrac{2d}{v}=\dfrac{2\cdot400\ ft}{1100\ \dfrac{ft}{s}}=0.73\ s."